OpenStudy (anonymous):

Okay i dont understand what they are asking me to do Which part of the quadratic formula tells you whether the quadratic equation can be solved by factoring? -b b2 - 4ac 2a Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

OpenStudy (hoblos):

b2 - 4ac for 2x2 + 7x + 3 = 0 b2 - 4ac= 7^2 - 4*3*2 = 25

OpenStudy (ash2326):

Discriminant tells us if a quadratic equation can be solved or not. \[D=b^2-4ac\] option b

OpenStudy (radar):

\[b ^{2}-4ac\] is termed the discriminant and if that value: is positive the equation has two real solutions is 0 the equation has one (repeated) solution is negative the equation has no real solutions (solutions are complex numbers)